OpenStudy (anonymous):

Choose whether the following is always sometimes or never true. x^3 = 3x

OpenStudy (istim):

Why not sub in dummy values and check.

OpenStudy (sasogeek):

it's sometimes true... and here's my reason. when x=0 0^3=3(0) 0=0 but it is not always true because if it was always true, then when x=1 1^3=3(1) 1=3 but 1 is not equal to 3 hence it is not always true, but can sometimes be true because it works for x=0 :)

OpenStudy (sasogeek):

so the final answer is, it is sometimes true :)